# Continual Adversarial Defense

### 15 Dec 2023 v1

@@ 6 Conclusion
The operational environment for defense systems is inherently dynamic, and it is unrealistic to expect a defense method to address all types of attacks. As a response to this situation, we propose the first continual adversarial defense (CAD) framework, designed to dynamically adapt to a wide array of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles including few-shot feedback and memory-efficient adaptation, and CAD demonstrates strong performance when operating under these principles. Through experiments conducted on CIFAR-10 and ImageNet-100, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
7 Acknowledgements
This work was supported in part by the Natural Science Foundation of China under Grant 61972169,62372203 and 62302186, in part by the National key research and development program of China(2022YFB2601802), in part by the Major Scientific and Technological Project of Hubei Province (2022BAA046, 2022BAA042), in part by the Knowledge Innovation Program of Wuhan-Basic Research, in part by China Postdoctoral Science Foundation 2022M711251.
Supplementary Material
@@ A Weight Estimator Base on Self-Perturbation
In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation to train a weight estimator to ensemble the target model and the defense model. We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples. Then, we train a 4-layer ConvNet as the weight estimator $f_{p}$ in a binary classification manner. In this section, we introduce the details of self-perturbation and model ensemble.

### 13 Mar 2024 v2

@@ 6 Conclusion
The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks. In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework. CAD is engineered to collect few-shot defense feedback from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles including few-shot feedback and memory-efficient adaptation, and CAD demonstrates strong performance when operating under these principles. Through experiments conducted on CIFAR-10 and ImageNet-100, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
@@ A Weight Estimator Base on Self-Perturbation
In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation [37] to train a weight estimator to ensemble the target model and the defense model. We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples. Then, we train a 4-layer ConvNet as the weight estimator $f_{p}$ in a binary classification manner. In this section, we introduce the details of self-perturbation and model ensemble.
@@ B Premium-CAD
C Implementation Details
C.1 Adversarial Attack
In this work, all of the adversarial attacks are implemented using TorchAttacks [15]. The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44]. Iterative steps and step size of iterative attacks are set to $10$ and $2/255$ respectively. The number of perturbed pixels of OnePixel [31] is set to $50$. The other parameters are the default ones in TorchAttacks.
C.2 Defense Baselines
The adversarial training baselines are implemented using RobustBench [3], while purification baselines are implemented using open-source code and model checkpoints. CL baselines are implemented using PyCIL [48]. The backbone models of both adversarial training methods and CL methods are standardized as WideResNet-28-10 for CIFAR-10 and ResNet-50 for ImageNet-100. CL models are trained under the same settings as CAD. Specifically, the training epochs and the amount of feedback per class for training CL models are set to 100 and K=10, respectively.